OpenStudy (anonymous):

a square has a side length of 95. what is the length of the diagonal of the square? Express your answer in simplest radical form.

OpenStudy (anonymous):

For equare; Diagonal Length² = Side Lenght² + Side Length² Diagonal Length² = 2 * (Side Lenght²) Therefore; Diagonal Length² = 2 * (95²) Diagonal Length = √(2 * (95²)) Diagonal Length = 95 * √(2) So the answer is 95*√(2) this is the same as √(18050)

OpenStudy (campbell_st):

use pythagoras' theorem 95^ + 95^2 = d^2

OpenStudy (anonymous):

\[a ^{2} + b^{2} = c^{2}\] \[95^2 + 95^2 = c^2\] \[c^2= 18050\] Find c